The graph of the function f(x)=mx+b contains the points (5,1) and (-10,-11). Find m and b.
Answer by mananth(16949)

f(x)=mx+b contains the points (5,1) and (-10,-11). Find m and b.
y= mx+b
plug the points and get a linear system
1=5m+b.............1
-11=-10m+b.........2
multiply (1) by 2
and add it to (2)
2=10m+2b
-11=-10m+b
-9=3b
b=-3
plug value of b in (1)
m=4/5
m.ananth@hotmail.ca